OpenStudy (anonymous):

Given the function f(x) = 3x, find the value of f-1(81)

OpenStudy (zzr0ck3r):

you need to solve \(3x=81\)

OpenStudy (zzr0ck3r):

\(f(a) = b \iff f^{-1}(b) = a\)

OpenStudy (dan815):

dooo itt

OpenStudy (anonymous):

To find the inverse: Replace f(x) with y Switch x's and y's, so put x where y is and x where y is. Solve for y Replace y with f^-1(x)

OpenStudy (dan815):

if 3x=y and x=1/3y and finv tells you what x you put in given a resultant y so x=1/3y is your f^-1 y=81 x=1/3*81=27

OpenStudy (dan815):

gimme them medals

OpenStudy (anonymous):

You aint pellet

OpenStudy (dan815):

LOL

OpenStudy (zzr0ck3r):

or just solve 3x=81